OpenStudy (loser66):

What is the branch cut? Problem: Prove that there is no branch of the logarithm defined on C - {0} Please, help

OpenStudy (anonymous):

there is no branch defined on all of \(\mathbb{C}\setminus \{0\}\) because of the discontinuity along the negative real axis, where the argument 'jumps' \(2\pi i\)

OpenStudy (loser66):

We can't just say that on the proof, right? we need an algebraic proof for this, right?

OpenStudy (loser66):

By contradiction, suppose there is a branch on S = C -{0}, then \(\forall z \in S\) log z = log|z| + i arg z And we need prove it leads to something wrong. That is what I think but not know how to put it in need yet.

OpenStudy (loser66):

Thanks for the link @thomas5267

OpenStudy (thomas5267):

I think the problem is that you have to choose a branch of \(arg(z)\) or else it won't work.

OpenStudy (loser66):

We cannot choose a case "it won't work" to prove the whole thing "won't work". I think.

OpenStudy (anonymous):

\(\log(z) = \log |z|+ i \arg(z)\) is the principal branch; all other branches differ by a multiple of \(2\pi i\). show that they are not continuous let alone analytic at a point the negative real axis (say, \(z=-1\))

OpenStudy (loser66):

So, we can prove the principal branch doesn't work on S, right?

OpenStudy (thomas5267):

Since a complex number can have \(\operatorname{Arg}(z)+2\pi k\) as argument and it is still valid, in order to make the logarithm a function you have to choose a branch of argument. No branch of argument is continuous.

OpenStudy (loser66):

like what oldrin says, let z =-1, then Log z = log|z|+iarg z= i(pi + 2kpi)

OpenStudy (loser66):

The function log z = ln |z| + i arg z is continuous on \(\mathbb C\setminus \{re^{i\alpha}, r\geq 0\}\)

OpenStudy (thomas5267):

Put it in another way, \(e^{i\theta}\) is surjective and not injective. \(e^{ix}=e^{iy}\not\implies x=y\).